# LSEAttention is All You Need for Time Series Forecasting

### v3

@@ Abstract
Transformer-based architectures have achieved remarkable success in natural language processing and computer vision. However, their performance in multivariate long-term forecasting often falls short compared to simpler linear baselines. Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain.To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting. We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models. Notably, LATST manage to achieve competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
@@ 1 Introduction
Multivariate time series forecasting is crucial in diverse domains such as finance, healthcare, and environmental monitoring, where the goal is to predict future values based on historical data. This task is particularly challenging in long-term forecasting, as it necessitates models that can effectively capture feature correlations and long-term dependencies across multiple time series. Recent research has increasingly focused on applying Transformer architectures to time-series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms. However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Zenget al. [1], raising concerns about the efficacy of Transformers in this context, particularly given their suboptimal performance and inherent complexity.
@@ 1.1 Transformer-based Long-term Time Series Forecasting
The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model which established its potential as a universal backbone for various tasks across multiple modalities.[2] The attention mechanism at the core of this architecture is formulated as follows:
@@
Recent studies have tailored Transformer models specifically for long-term time-series forecasting. For example, Informer introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals.[3] Similarly, Autoformer employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.[4]
FEDformer incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.[5] Instead of concentrating solely on point-wise attention, PatchTST emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.[6]
The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert in their SAMformer model. SAMformer integrates Reversible Instance Normalization (RevIN) to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.et al. [8, 7]This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
@@ 1.3 Contributions
  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the attention for time-series data.
  • Introduction of the LATST (LESAttention Time-Series Transformer) a Transformer with LSEAttention replacement that attempts to solve the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
  • Empirical validation of our approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art Transformers.
@@ 2.2 Motivation
Transformers have emerged as a powerful architecture for capturing temporal associations in sequential data, primarily due to their reliance on point-wise attention mechanisms. However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly identical matrix. This identity diminishes the model’s ability to differentiate patterns, thereby hindering generalization. An equally critical yet less-explored issue is entropy collapse, particularly prevalent in the application of Transformers to time-series forecasting. In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation.[1] A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
In the attention mechanism Eq. 1, attention scores are computed as the dot product of query and key vectors. This operation tends to disproportionately emphasize abrupt changes or noise by assigning them larger values. Subsequently, when these scores are passed through the softmax function, defined as:
@@
The exponential computation introduces significant numerical instability. Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow.[10] This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation. Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output.
This disproportionate weighting undermines the diversity of the attention mechanism, reducing its capacity to effectively model complex temporal relationships. As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data which is the inherent characteristic of time-series data. This behavior culminates in entropy collapse, where the attention distribution becomes overly concentrated on a few elements, thereby limiting the model’s representational power. In other words, trend and periodicity might be barely considered.[1] Addressing this issue necessitates exploring alternatives to the conventional softmax function.
@@
To address these challenges, we propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eq. 4 trick proposed by Blanchard et al. [10] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al. [11] to form a gated attention . The LSE trick alleviates entropy collapse issue arising from numerical instability through subtraction of max value. Starting from LSE where the softmax can be represented with LSE in this way:
@@
where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing the numerical stability.
@@
$y=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)=\log\left(a\sum_{i=1}^{n}e^{x_{i}}-a\right)$ (6)
where $a$ is a constant and maximum value utilized ensuring that the largest exponent becomes zero and the others become non-positive. As a result, all exponential terms are bounded between 0 and 1 like normalization for numerical stabilization purpose in practice. This can be simplified to:
@@
To enhance the model’s capabitlity, the GELU is applied to introduce the non-linearity, then the equation is modfied as:
@@
$g_{i}=\exp(x_{i}-y)$ (9)
@@ 2.3 LATST: Overall Structure
The architecture replaces the traditional temporal self-attention mechanism, with the LSEAttention module. To assess the efficacy of the LSEAttention component, we conducted an ablation study comparing LATST with the Gaussian Error linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified linear Unit (ReLU). The details of this study can be found in Ablation Study of Sec. 3.
Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing adaptive learning while maintaining the robustness of attention mechanisms. This design facilitates effective modeling of temporal dependencies and enhances performance in time series forecasting tasks.
Figure 1: Overall structure of LATST.
@@ 3.1 Dataset
Evaluation Metric: Consistent with previous studies, we employ the widely used evaluation metric, mean squared error (MSE). This metric has been a standard choice in earlier works and is defined as follows:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(Y_{i}-\hat{Y}_{i})^{2},$ (10)
where $N$ is the total number of samples, $Y_{i}$ is the $i$-th prediction, and $\hat{Y}_{i}$ represents the corresponding ground truth value.
@@ 3.2 Baseline
Initial learning rate 0.5 to 0.00005
@@
Table 3: Parameters comparison of LATST and SAMformer models for various datasets and prediction lengths
@@
Electricity 251K 263K 282K 331K 82K 131K 205K 401K
Traffic 251K 263K 282K 331K 82K 131K 205K 401K
Weather 51K 54K 58K 72K 82K 131K 205K 401K
ETTH1 29K 29K 58K 331K 82K 131K 205K 401K
ETTH2 25K 27K 63K 331K 82K 131K 205K 401K
ETTM1 109K 115K 125K 331K 82K 131K 205K 401K
ETTM2 109K 115K 125K 150K 82K 131K 205K 401K
Exchange 51K 115K 199K 236K 82K 131K 205K 401K
@@ 3.3 Performance Analysis
LATST demonstrates exceptional performance, achieving the lowest mean squared error (MSE) scores in 21 out of 32 scenarios across eight diverse datasets. This remarkable consistency underscores the model’s robustness and adaptability, effectively addressing a wide range of time-series forecasting tasks. Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, demonstrating its superior ability to manage complex temporal dependencies.
When comparing LATST directly with SAMformer, the model shows its superiority by outperforming SAMformer on 6 out of 8 datasets with overall improvement by 5.12% . Moreover, LATST surpasses the traditional Transformer model by a notable margin of 19.75%, which emphasizes its enhanced predictive capabilities. Key innovations in LATST, such as the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute significantly to this performance enhancement. These modifications enable LATST to capture long-range dependencies and temporal patterns more effectively, setting it apart from its counterparts.
The detailed performance metrics for each dataset and prediction length are summarized in Tab. 2. Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variants. This phenomenon is closely related to the number of variants and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions. A more in-depth analysis can be found in Tab. 3.
In comparison with DLinear, LATST outperforms the model on 5 out of 8 datasets, despite DLinear having 139.7K parameters. LATST, with fewer parameters in most scenarios, demonstrates that its performance is not solely dependent on model size, but rather on the efficiency and effectiveness of its architecture.
Consistency Across Datasets: LATST’s ability to achieve the best MSE scores in a significant number of scenarios further reinforces its robustness and generalizability across diverse forecasting challenges. This performance stability across different datasets suggests that LATST effectively learns and captures the underlying temporal patterns, allowing it to adapt well to varying data distributions and dynamics. For instance, in the Traffic dataset, LATST consistently maintains lower MSE values even at longer prediction horizons, illustrating its resilience against overfitting and its capacity to generalize beyond training conditions.
Table 4: Performance of LATST with different activation functions for prediction length of 96.
Dataset With PReLU With ReLU
ETTh2 0.290 0.292
ETTm1 0.295 0.305
ETTm2 0.176 0.191
Exchange Rate 0.098 0.115
@@ 3.4 Ablation Study
In this section, we present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework. By systematically removing or modifying key elements, such as the GeLU and the PReLU, their contributions are isolated to the overall performance of the model. For consistency, we fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
Table 5: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
Dataset Without GeLU Transformer
ETTh2 0.312 0.396
ETTm1 0.333 0.384
ETTm2 0.183 0.200
Exchange Rate 0.105 0.292
Preliminary results from the ablation study, as shown in LABEL:tab:ab_relu, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets. This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
@@
As illustrated in LABEL:tab:ab_gelu, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration. This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function. The improvement achieved through the Log-Sum-Exp trick is attributed to its ability to stabilize gradient computations during training, thereby facilitating the learning of more complex patterns.
@@ 4 Conclusion and Future Work
In this paper, we present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting. Our empirical results confirm the effectiveness and robustness of the proposed LATST approach inty of forecasting contexts. Although we do not exhaustively explore all potential causes of entropy collapse, we introduce a combination of novel modules specifically combined to address this issue.

### v4

# LATST: Are Transformers Necessarily Complex for Time-Series Forecasting

@@ Abstract
Transformer-based architectures have achieved remarkable success in natural language processing and computer vision. However, their performance in multivariate long-term forecasting often falls short compared to simpler linear baselines. Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain.To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting. We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models. Notably, LATST achieves competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
@@ 1 Introduction
Multivariate time series forecasting is crucial in diverse domains such as finance, healthcare, and environmental monitoring, where the goal is to predict future values based on historical data. This task is particularly challenging in long-term forecasting, as it necessitates models that can effectively capture feature correlations and long-term dependencies across multiple time series. Recent research has increasingly focused on applying Transformer architectures to time-series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms. However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Ni et al. [1], raising concerns about the efficacy of Transformers in this context, particularly given their suboptimal performance and inherent complexity.
@@ 1.1 Transformer-based Long-term Time Series Forecasting
The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model [3] which established its potential as a universal backbone for various tasks across multiple modalities. The attention mechanism at the core of this architecture is formulated as follows:
@@
Recent studies have tailored Transformer models specifically for long-term time-series forecasting. For example, Informer [4] introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals. Similarly, Autoformer [5] employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.
FEDformer [6] incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences. Instead of concentrating solely on point-wise attention, PatchTST [7] emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.
The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert et al. [8] in their SAMformer model. SAMformer integrates Reversible Instance Normalization (RevIN) [9] to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability. This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
@@ 1.3 Contributions
  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the softmax operation.
  • Introduction of the LSEAttention module, a variation of traditional multi-head attention that addresses the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
  • Empirical validation of my approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art models.
@@ 2.2 Motivation
Transformers have emerged as a powerful architecture for capturing temporal associations in sequential data, primarily due to their reliance on point-wise attention mechanisms. However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly uniform distributions across diverse inputs. This uniformity diminishes the model’s ability to differentiate patterns, thereby hindering generalization. An equally critical yet less-explored issue is entropy collapse, particularly prevalent in the application of Transformers to time-series forecasting. In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation [1]. A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
In the attention mechanism (Equation 1), attention scores are computed as the dot product of query and key vectors. This operation tends to disproportionately emphasize abrupt changes or noise by assigning them larger values. Subsequently, when these scores are passed through the softmax function, defined as:
@@
the exponential computation introduces significant numerical instability. Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow. This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation. Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output [11].
This disproportionate weighting undermines the diversity of the attention mechanism, reducing its capacity to effectively model complex temporal relationships. As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data. This behavior culminates in entropy collapse, where the attention distribution becomes overly concentrated on a few elements, thereby limiting the model’s representational power. In other words, trend and periodicity are barely considered. Addressing this issue necessitates exploring alternatives to the conventional softmax function.
@@
To address these challenges, We propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eqn4 trick proposed by Blanchard et al. [11] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al. [12]. The LSE trick alleviates entropy collapse issue arising from numerical instability through subtraction of max value. By starting from LSE where the softmax can be represented with LSE in this way:
@@
where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing numerical stability.
@@
$y=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)=\log\left(e^{a}\sum_{i=1}^{n}e^{x_{i}-a}\right)$ (6)
where $a$ is a constant and max value utilized for ensuring that the largest exponent becomes zero and the others become non-positive. As a result, all exponential terms are bounded between 0 and 1 like normalization in practice. This can be simplified to:
@@
To enhance the model’s capabitlity, the GELU is applied to indroduce the non-linearity, then the equation is modfied as:
@@
$g_{i}=\exp(x_{i}-y)=\frac{e^{x_{i}}}{e^{y}}$ (9)
$\frac{e^{x_{i}}}{e^{\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)}}=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=\text{softmax}(x_{i})$ (10)
@@ 2.3 LATST: Overall Structure
The architecture retains a traditional temporal self-attention mechanism, integral to the LSEAttention module. To assess the efficacy of the LSEAttention component, We conducted an ablation study comparing LATST with the Gaussian Error linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified linear Unit (ReLU). The details of this study can be found in Ablation Study of Sec. 3.
Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing for adaptive learning while maintaining the robustness of attention mechanisms. This design facilitates effective modeling of temporal dependencies and enhances performance in time series forecasting tasks.
Figure 1: Overall structure of LATST
@@ 3.1 Dataset
Evaluation Metric: Consistent with previous studies, We employ the widely used evaluation metric, mean squared error (MSE). where $N$ is the total number of samples, $Y_{i}$ is the $i$-th prediction, and $\hat{Y}_{i}$ represents the corresponding ground truth value.
@@ 3.2 Baseline
Initial learning rate 0.5 to 0.00001
@@
Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
Electricity Avg 0.163 0.181 0.209 0.209 - 0.166
Traffic Avg 0.415 0.425 0.457 0.620 0.428 0.434
Weather Avg 0.233 0.260 0.279 0.267 0.258 0.246
ETTh1 Avg 0.414 0.410 0.554 0.439 0.454 0.423
ETTh2 Avg 0.355 0.344 0.412 0.357 0.383 0.431
ETTm1 Avg 0.363 0.373 0.427 0.385 0.407 0.357
ETTm2 Avg 0.266 0.269 0.302 0.289 0.288 0.267
Exchange Avg 0.412 0.445 0.620 0.593 0.360 0.287
@@
Electricity Avg 251K 263K 282K 331K 82K 131K 205K 401K
Traffic Avg 251K 263K 282K 331K 82K 131K 205K 401K
Weather Avg 50.5K 53.7K 58.5K 71.7K 82K 131K 205K 401K
ETT Avg - - - - 82K 131K 205K 401K
Exchange Avg - 115K - 236K 82K 131K 205K 401K
@@ 3.3 Performance Analysis
LATST demonstrates superior performance by achieving the best mean squared error (MSE) scores in 21 out of 32 scenarios across the evaluated eight datasets. This remarkable consistency underscores the model’s robustness and adaptability in addressing diverse time series forecasting tasks. Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, indicating its efficacy in managing complex temporal dependencies.
When comparing LATST directly with SAMformer, LATST exhibits an overall improvement of 4.3% in MSE. Furthermore, LATST surpasses the Transformer model by a substantial margin of 19.11%, highlighting its enhanced predictive capabilities. The modifications introduced in LATST, including the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute to this performance enhancement, allowing the model to better capture long-range dependencies and temporal patterns.
The detailed performance metrics for each dataset and prediction length are summarized in Tab. 2. Notably, LATST consistently achieves the lowest MSE across multiple input lengths for datasets such as ETTh1, ETTh2, and Electricity. For instance, in the Electricity dataset, LATST records an MSE of 0.131 for a prediction length of 96, significantly outperforming the closest competitor, SAMformer, which achieves an MSE of 0.155. This trend is replicated across various prediction horizons, demonstrating the model’s ability to maintain stability and robustness under different forecasting conditions.
Consistency Across Datasets: The model’s ability to achieve the best MSE scores in a significant number of scenarios reinforces its robustness and generalizability to diverse forecasting challenges. The performance stability observed across different datasets suggests that LATST effectively learns and captures underlying temporal patterns, enabling it to adapt well to varying data distributions and dynamics. For example, in the Traffic dataset, LATST maintains lower MSE values even at longer prediction horizons, indicating its resilience against overfitting and its capacity to generalize beyond training conditions.
@@ 3.4 Ablation Study
In this section, We present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework. By systematically removing or modifying key elements, such as the GeLU and the PReLU, We isolate their contributions to the overall performance of the model. For consistency, We fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
Dataset Pred. Length With PReLU With ReLU
ETTh2 96 0.290 0.292
ETTm1 96 0.300 0.305
ETTm2 96 0.176 0.191
Exchange Rate 96 0.098 0.115
Table 3: Performance of LATST with different activation functions
Preliminary results from the ablation study, as shown in Tab. 3, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets. This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
@@
Dataset Pred. Length Without GeLU Transformer
ETTh2 96 0.312 0.396
ETTm1 96 0.333 0.384
ETTm2 96 0.183 0.200
Exchange Rate 96 0.105 0.292
Table 4: Performance of LATST without GeLU compared to Transformer
As illustrated in Tab. 4, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration. This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function. The improvement achieved through the Log-Sum-Exp trick is attributed to its ability to stabilize gradient computations during training, thereby facilitating the learning of more complex patterns.
@@ 4 Conclusion and Future Work
In this paper, we present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting. Our empirical results confirm the effectiveness and robustness of the proposed LATST approach inty of forecasting contexts. Although we do not exhaustively explore all potential causes of entropy collapse, We introduce a novel module specifically designed to address this issue.